OpenStudy (anonymous):

Find critical number of the function f(x) = 18 x \ln x

OpenStudy (diamondboy):

Hi :)

OpenStudy (diamondboy):

start by first finding the derivative then set it to 0

OpenStudy (anonymous):

I got 0 = 18x (1/x) + (lnx)(18). (its suppose to be multiply not divide sorry) (-lnx18x)/18 = x

OpenStudy (anonymous):

i would say \[18+18\ln(x)\] since \[\frac{18x}{x}=18\]

OpenStudy (anonymous):

how did you get 18 + 18ln(x)